OpenStudy (anonymous):

Slove the Following Multi-Step Equation -12+ -W/4=-6 (-W/4 is a fraction)

OpenStudy (texaschic101):

-12 + (-w/4) = -6 -- distribute through the parenthesis -12 - w/4 = -6 if you have an equation with fractions, and you want to get rid of the fractions, just multiply by the common denominator , which in this case, is 4. -12(4) - (w/4)(4) = -6(4) -48 - w = -24 -- add w to both sides -48 = -24 + w --- add 24 to both sides -48 + 24 = w - 24 = w check.. -12 + (-w/4) = -6 -12 - w/4 = -6 -12 - (-24/4) = -6 -12 + 6 = -6 -6 = -6 (correct) w = -24

OpenStudy (anonymous):

Thank you!